OpenStudy (jjuden):

which expression is equivalent to (sqrt x-5)^7 a(x-5)^1/7 b(x-5)^7/1 c(x-5)^7/2 d(x-5)^2/7

OpenStudy (jjuden):

\[(\sqrt{x-5})^7\]

OpenStudy (bibby):

\(\huge \sqrt{a}=a^\frac{1}{2} \\\huge (\sqrt{x-5})^7=((x-5)^\frac{1}{2})^7\) now recall that when you raise one power to another, you multiply the exponents. \(\huge (a^b)^c=a^{(b*c)}\)

OpenStudy (bibby):

so c

OpenStudy (jjuden):

thats what i got and its right lol thank you :)